OpenStudy (anonymous):

What’s the equation of a line that passes through points (0, -1) and (2, 3)? A. y = 2x − 1 B. y = 2x + 1 C. x = 2y − 1 D. x = 2y + 1

OpenStudy (anonymous):

If anyone views this could you also tell me how you got the answer so I know in the future

OpenStudy (anonymous):

y=2x-1 -1=2*0-1 -1=-1 3=2*2-1 3=4-1 3=3 it is the first one

OpenStudy (anonymous):

Okay thanks @sodukogirl13